### v3

@@ 3 Discussion
Theories of concepts have identified various properties that conceptual representations must satisfy, highlighted by the symbolism-connectionism divide. Our results show that LLM-derived conceptual representations successfully reconcile the seemingly competing properties, integrating the definitions, relations and structures emphasized by the symbolic approaches with the graded and continuous nature of neural networks. These representations were structured in a way that their relationships supported straightforward computations of human-like similarities, categories (Fig. 4, Supplementary Information SI 1.4) and gradient distinctions (Fig. 5). Previous work has employed distributed word embeddings to approximate these aspects of human concept usage, revealing preliminary correspondence [47, 54, 49]. However, these embeddings have exhibited inconsistent alignment with human similarity judgments across various datasets [27]. Our results align with previous findings, showing that word embeddings primarily reflect association or relatedness but fail to capture genuine similarity (Fig. 4a–b, Supplementary Information SI 1.3). In contrast, LLM-derived conceptual representations demonstrated distinctive strengths. They excelled in modeling human similarity judgments, including those based on images, suggesting that they advantageously capture conceptual information independent of stimulus modality. These representations also supported flexible, context-dependent reorganization to reflect gradient scales along different features, surpassing word embeddings in both breadth and depth. Importantly, they exhibited superior alignment with neural activity patterns in the human brain (Fig. 7). Collectively, these findings underscore the potential of LLM-derived conceptual representations to advance our understanding of human concepts and to bridge the long-standing divide between symbolic and connectionist approaches.
Vector-based models have been argued to plausibly capture the neural activations underlying actual cognitive processes in the brain [16, 8]. The mapping between LLM-derived conceptual representations and brain activity patterns suggests that they provide a plausible basis for analyzing neural representations of concepts (Fig. 7). Compared to word embeddings and representations derived from human similarity judgments, the LLM-derived representations exhibited notable advantages, especially in accounting for neural activity patterns in high-level visual regions (Fig. 7d–i) that are associated with abstract semantic information [50, 51, 53]. However, they fell short in capturing certain aspects of behaviorally relevant, low-level visual information (Fig. 7d–f). To understand this discrepancy, we regressed the similarity representations onto the LLM-derived representations and found the most pronounced gaps in factors related to color, followed by texture and shape, which are primarily perceptual properties (Supplementary Information SI 1.6, Fig. S18). This result aligns with previous research showing that blind individuals primarily acquire such properties through inference and differ significantly from sighted people in their knowledge of color [37, 66]. Such information, therefore, may be inefficient to learn from language or even absent altogether. Despite these limitations, the substantial alignment between LLM-derived representations and neural responses to visual stimuli adds to the evidence that language prediction can orient models towards a shared representational structure of reality, even in the absence of real-world grounding [36, 64]. While prior work has hinted at this convergence [67], our findings explicitly characterize the conceptual structure emerging solely from language, setting the stage for future research on the alignment and interaction between linguistic and visual systems.
Our work provides a foundational framework for exploring the emergence of conceptual representations from language prediction. We demonstrated that deep neural networks, trained for next-token prediction at sufficient scale and with extensive high-quality training data, can develop abstract conceptual representations that converge toward a shared, context-independent relational structure, enabling generalization. Despite the lack of real-world reference, these representations reconcile essential properties of human concepts and plausibly reflect neural representations in the brain associated with visually grounded concepts. These findings position LLMs as powerful tools for advancing theories of concepts. However, unlike human cognition that relies on concepts for understanding and reasoning, LLMs operate at the token level and do not necessarily utilize the conceptual representations during typical text generation. This discrepancy may lead to the observed limitations of LLMs in certain aspects of concept usage, such as compositionality [27] and reasoning [4, 21].
@@
Future work could aim to build better models of human cognition by incorporating more cognitively plausible incentives [68] such as systematic generalization [69] and reasoning [70, 71, 72], while explicitly guiding models to leverage conceptual representations, rather than linguistic forms, to generalize across tasks. Recent progress in steering LLMs to operate within their representation spaces have shown promise in enhancing both language generation and reasoning [73, 74]. Efforts in this direction could narrow the gaps between LLMs and human conceptual abilities and help elucidate their current limitations in compositionality [27], reasoning [21, 4] and over-sensitivity to minor contextual shifts [21, 75]. Moreover, the LLM-derived conceptual representations could be enriched with information from diverse sources, like vision [76, 63], to better align with human cognition [65] and foster human-machine collaboration [77, 78]. Finally, incorporating brain data beyond the visual domain would offer a richer understanding of the neural underpinnings of conceptual representations in the human mind. Despite the current limitations in models and data, the emergence of human-like conceptual representations within LLMs marks a critical step towards resolving enduring questions in the science of human concepts. This progress opens new avenues for bridging the gaps between human and machine intelligence, offering valuable insights for both cognitive science and artificial intelligence.
4 Methods
@@ 4.1 Large language models used in our experiments
This paper focuses on base models, i.e., LLMs pretrained solely for next-token prediction without fine-tuning or reinforcement learning. Our experiments exclusively utilized open-source LLMs, as their hidden representations are necessary for our analysis. We primarily conducted our experiments on the LLaMA 3 models, including LLaMA3-70B and LLaMA3-8B, both trained on over 15 trillion tokens [79]. Another 11 series of Transformer-based decoder-only LLMs were also employed for experiments on the reverse dictionary task and the convergence of LLM representations. These included (1) Falcon [80], (2) Gemma [81], (3) LLaMA 1 [82], (4) LLaMA 2 [83], (5) Mistral [84, 85], (6) MAP-Neo [86], (7) OLMo [87], (8) OPT [88], (9) Phi [89], (10) Pythia [90], and (11) Qwen [91, 92]. Additional details about the models’ names, scales, training data and sources can be found in Table S1 and Table S2. These models vary in architecture, scale, and pretraining data, enabling explorative analyses of how these factors might impact the conceptual representations and organization within LLMs.
@@ 4.2 Details on deriving conceptual representations
We used data from the THINGS database [42] to probe LLMs’ conceptual representations via the reverse dictionary task. The dataset includes 1,854 concrete and nameable object concepts, paired with their WordNet synset IDs, definitional descriptions, several linked images and category membership labeled by humans. The concepts and images were selected to be representative of everyday objects commonly used in American English, providing a useful testbed for analyzing model representations. To assess the generality of our results, we extended our analysis to a broader range of descriptions and concepts. Specifically, (1) we tested LLMs’ generalizability across different word classes (nouns, verbs, adjectives and adverbs), age-of-acquisition, and degrees of concreteness using a broader set of 21,402 concepts. The concepts were selected from the intersection of the three datasets: age-of-acquisition [93], concreteness [94] and Wordnet [95]. (2) We evaluated the consistency of LLMs’ predictions using three additional distinct definitional descriptions for each of the 1,854 THINGS concepts, which were generated by GPT-3.5 and manually checked for diversity (see Supplementary Information SI 2.1 for details). (3) We examined LLMs’ sensitivity to linguistic structure by introducing varying degrees of word order permutations to the query descriptions. We shuffled $30\%$, $60\%$ and $100\%$ of the words in the query descriptions from the THINGS database and reinserted them into the original text.
@@
To probe the interrelatedness among concepts within LLMs, we examined how modifying the description-word pairings in context affects model inferences. For each target concept, we paired its description with a proxy symbol and combined it with $N-1$ correct description-word pairs from other concepts. We then queried the LLM using the same description. Model performance was evaluated based on how often the LLM replicated the proxy symbol or generated the correct word for the target concept. We tested four types of proxy symbols: (1) a random uppercase English letter, excluding “A” and “I” to avoid potential semantic associations; (2) a randomly generated lowercase letter string, with length sampled from a shifted Poisson distribution ($\textrm{mean}=6.94$, variance = 5.80) to approximate typical English word lengths [97]; (3) a random word selected from the THINGS database, distinct from the target concept and absent from the given context; and (4) the correct word for the target concept, used as a baseline.
@@ 4.3 Structural analysis of representation spaces
We used representational similarity analysis (RSA) [44] to measure the alignment between conceptual representations derived under different conditions (i.e., varying contexts or models), which is non-parametric and has been widely adopted to measure topological alignment between representation spaces [98]. Let $X\in\mathbb{R}^{m\times d_{1}}$ and $Y\in\mathbb{R}^{m\times d_{2}}$ denote two sets of conceptual representations for $m$ concepts with dimensionality of $d_{1}$ and $d_{2}$, respectively. Each space is characterized by its relational structure, represented by a (dis)similarity matrix $M\in\mathbb{R}^{m\times m}$, where the entry $M_{i,j}$ denotes the (dis)similarity between the representations of the $i^{\textrm{th}}$ and $j^{\textrm{th}}$ concepts in the corresponding space. The alignment can then be calculated as the Spearman’s rank correlation between the upper (or lower) diagonal portion of the two matrices $M_{X}$ and $M_{Y}$, yielding values ranging from $-1$ to $1$.
An appropriate similarity function is needed for the (dis)similarity matrix to characterize the relational structure of the representation space. While cosine similarity has been widely adopted since the advent of static distributed word embeddings, it may be suboptimal for capturing non-linear relationships and sensitive to outliers [99]. We thus employed two different metrics including the cosine similarity and Spearman’s rank correlation. Comparison with human behavioral data suggests that the cosine similarity captures semantic similarity (e.g., cups and mugs) while Spearman’s rank correlation reflects both similarity and association (e.g., cups and coffee) (Supplementary Information SI 1.3, Fig. S10a–b). We primarily used Spearman’s rank correlation as the similarity function, with results from alternative metrics provided in Supplementary Information SI 1.2.
Additionally, we employed the parallelism score (PS) [58] to evaluate if the differences between conceptual representations are preserved across conditions, which have been thought to reflect relations in the representation space [100, 49]. The PS for each concept pair, among the total $m$ concepts, is computed as the cosine similarity between the vector offsets from one concept to the other. We reported the average PS for two representation spaces (Supplementary Information SI 1.2, Fig. S9c–d).
For comparison across different LLMs, we also visualized these models using t-SNE with a perplexity of 30 and 1000 iterations, where distances were computed as $1-\textrm{alignment}$.
@@ 4.4 Characterization of model complexity
For a general characterization of the complexity of the 67 LLMs used in our experiments, we represented each by its number of parameters and amount of training data (i.e., the number of tokens used during training)—factors identified as crucial for determining model quality [101, 102]. A principal component analysis (PCA) was then applied to these factors. The first principal component, accounting for $70.67\%$ of the total variance, was used to represent the complexity of each model. In terms of the Mistral models, where information about their pretraining data was not accessible, we estimated it based on the data volume of comparable models released around the same time (Table S1).
@@ 4.5 Prediction of human similarity judgments
To evaluate how closely the relationships between LLM-derived representations align with psychological measures of similarity, we first compared them with human similarity ratings for concept pairs from SimLex-999 [47]. The dataset explicitly distinguishes semantic similarity from association or relatedness and contains human ratings for 999 concept pairs spanning 1,028 concepts. The concepts cover three word classes including nouns, verbs and adjectives. We used description-word pairs from the WordNet data to provide contextual demonstrations to the LLMs, thereby deriving conceptual representations. Model performance was assessed through Spearman’s rank correlation between the similarity scores derived from the LLM representations and the human ratings.
We further employed the odd-one-out similarity judgments from the THINGS-data collection [40] to validate the effectiveness of LLM representations in handling the computation of similarities. The dataset consists of triplets sampled from the 1,854 concepts in THINGS. We presented LLMs with contextual demonstrations randomly sampled from THINGS to obtain their conceptual representations. For each triplet $\left(i,j,k\right)$, we took the corresponding conceptual representations $\left(\mathbf{h}_{i},\mathbf{h}_{j},\mathbf{h}_{k}\right)$ and calculated pairwise similarity. The one outside the most similar pair was identified as the odd-one-out and compared to human judgments. We evaluated model performance mainly with a set of 1,000 triplets from THINGS, each with multiple responses collected from different participants. We compared the LLMs’ judgments with the majority of human choices, which provides a reliable estimation of the alignment between LLMs and humans.
@@ 4.6 Categorization
For the categorization experiment, we employed high-level human-labeled natural categories in THINGS [42]. We removed subcategories of other categories, concepts belonging to multiple categories and categories with fewer than ten concepts [48]. This resulted in 18 out of 27 categories, including animal, body part, clothing, container, electronic device, food, furniture, home decor, medical equipment, musical instrument, office supply, part of car, plant, sports equipment, tool, toy, vehicle and weapon. These categories comprise 1,112 concepts.
We employed three methods to evaluate the extent to which category membership can be inferred from LLM-derived conceptual representations, with the former two corresponding to prototype and exemplar models and the third explicitly examining the relationships between high-level categories and their associated concepts. Specifically, for the prototype model, we used a cross-validated nearest-centroid classifier, performing categorization by iteratively leaving each concept out. In each iteration, we calculated the centroid for each category by averaging the representations of the remaining concepts. Categorization was then based on the similarity between the left-out concept and each category centroid. In the exemplar model, categorization was carried out using a nearest-neighbour decision rule [103], where each concept was classified based on the category membership of its closest neighbours among all other concepts (Supplementary Information SI 2.2). Our final approach involved constructing a representation for each of the 18 categories using the same $N$ demonstrations employed for the 1,854 concepts. Here, we directly used the category names as query descriptions, categorizing concepts by their similarity to each category representation.
We combined t-SNE with multidimensional scaling (MDS) to visualize the representations in two dimensions, thereby preserving the global structure while better capturing local similarities. The representations were first reduced to 64 dimensions using MDS, with distances calculated as $1-\textrm{similarity}$ and then visualized using t-SNE with a perplexity of 30 and 1000 iterations.

### v4

@@ 3 Discussion
Theories of concepts have identified various properties that conceptual representations must satisfy, highlighted by the symbolism–connectionism divide. Our results show that LLM-derived conceptual representations successfully reconcile the seemingly competing properties, integrating the definitions, relations and structures emphasized by the symbolic approaches with the graded and continuous nature of neural networks. These representations were structured in a way that their relationships supported straightforward computations of human-like similarities, categories (Fig. 4; SI Appendix, section SI 1.E) and gradient distinctions (Fig. 5). Previous work has employed distributed word embeddings to approximate these aspects of human concept usage, revealing preliminary correspondence [49, 55, 52]. However, these embeddings have exhibited inconsistent alignment with human similarity judgments across various datasets [26]. Our results align with previous findings, showing that word embeddings primarily reflect association or relatedness but fail to capture genuine similarity (Fig. 4 AB; SI Appendix, section SI 1.C). In contrast, LLM-derived conceptual representations demonstrated distinctive strengths. They performed well in modeling human similarity judgments, including those based on images, suggesting that they advantageously capture conceptual information extending beyond specific stimulus modalities. They also supported flexible, context-dependent reorganization to reflect gradient scales along different features, surpassing word embeddings in both breadth and depth. Importantly, these representations exhibited superior alignment with brain activity patterns, especially in high-level visual areas associated with abstract semantic information (Fig. 7 DI), supporting the view that vector-based models offer a plausible framework for capturing the neural activations underlying cognitive processes [15, 8]. Collectively, these findings underscore the potential of LLM-derived conceptual representations to reconcile symbolic and connectionist approaches and to illuminate the representational basis of human concepts.
Despite this promise, our results also reveal notable divergences from human behavioral and neural patterns. LLM-derived conceptual representations fell below the noise ceiling in predicting human similarity judgments and accounted for less of the variance in low-level visual regions than embeddings derived from human similarity judgments. To probe this discrepancy, we regressed the human-derived similarity embedding onto the LLM-derived representations and found the most pronounced gaps in dimensions related to color, followed by texture and shape (SI Appendix, section SI 1.H; Fig. S12B). These gaps closely mirrored the error profile of the LLM-derived representations in behavioral predictions and mapped onto neural activity in low-level visual areas (SI Appendix, section SI 1.H; Figs. S12, S24), highlighting a key limitation in capturing visually grounded perceptual properties. These findings parallel previous research showing that congenitally blind individuals acquire these properties primarily through inference and differ markedly from sighted people in their knowledge of color [36, 68]. Such information, therefore, may be inefficient to learn from language alone, or even entirely absent.
Our work provides a foundational framework for exploring emergent conceptual representations from language prediction, positioning LLMs as powerful tools for advancing theories of concepts. While our approach to deriving conceptual representations from LLMs exhibits promising generalizability across datasets, tasks, and models, the human conceptual repertoire is far richer and more nuanced than what we have examined. Extending this approach to more diverse domains—such as abstract concepts, compositional structures, and cross-linguistic variation—and systematically investigating the factors that drive convergence and divergence between LLM and human representations [69] will be essential for fully characterizing model competence and limitations. Moreover, although the structure of the conceptual representations correlates strongly with downstream performance, LLMs operate at the token level and do not necessarily draw on these representations during typical text generation. This disconnect may underlie their observed limitations in certain aspects of concept usage, including compositionality [26] and reasoning [4, 20].
@@
Future work could aim to build better models of human cognition by incorporating more cognitively plausible incentives [70] such as systematic generalization [71] and reasoning [72], while explicitly guiding models to leverage conceptual representations, rather than linguistic forms, to generalize across tasks. Recent progress in steering LLMs to operate within their representation spaces has shown promise for enhancing language generation and reasoning [73, 74]. Efforts in this direction could narrow the gaps between LLMs and human conceptual abilities and help elucidate their current limitations in compositionality [26], reasoning [20, 4] and over-sensitivity to minor contextual shifts [20, 75]. Moreover, the LLM-derived conceptual representations could be enriched with information from diverse sources, like vision [76, 63], to better align with human cognition [67] and foster human-machine collaboration [77, 78]. Finally, incorporating brain data beyond the visual domain would offer a richer understanding of the neural underpinnings of conceptual representations in the human mind. Despite current limitations in models and data, the emergence of human-like conceptual representations within LLMs marks a critical step toward resolving enduring questions in the science of human concepts. This progress opens new avenues for bridging the gaps between human and machine intelligence, offering valuable insights for cognitive science and artificial intelligence.
4 Materials and Methods
@@ 4.1 Large language models used in our experiments
This paper focuses on base models—LLMs pretrained solely for next-token prediction without additional fine-tuning or reinforcement learning. We used only open-source LLMs, as their hidden representations are necessary for our experiments. Primary analysis were conducted on LLaMA 3 models, including LLaMA3-70B and LLaMA3-8B, each pretrained on over 15 trillion tokens [79]. Another 65 Transformer-based decoder-only LLMs from various model series were also employed for experiments on the reverse dictionary task and representational convergence. Further details regarding model names, scales, training data and sources can be found in SI Appendix, Tables S1S2. These models vary in architecture, scale, and pretraining data, enabling exploratory analyses of how these factors might shape the conceptual representations within LLMs.
@@ 4.2 Details on deriving conceptual representations
We used data from the THINGS database [40] to probe LLMs’ conceptual representations via the reverse dictionary task. The dataset includes 1,854 concrete and nameable object concepts, paired with their WordNet synset IDs, definitional descriptions, several linked images and human-labeled category membership. The concepts and images were selected to be representative of everyday objects commonly used in American English, providing a useful testbed for analyzing model representations. The dataset is available at https://osf.io/jum2f/. To assess the generality of our results, we extended our analysis to a broader range of descriptions and concepts. Specifically, (1) we tested LLMs’ generalizability across a broader set of 21,402 concepts, spanning different word classes (nouns, verbs, adjectives and adverbs), age-of-acquisition [80], and degrees of concreteness [81], all intersecting with Wordnet [82]. (2) We evaluated LLMs’ prediction consistency using three additional distinct definitional descriptions for each THINGS concept, which were generated by GPT-3.5 and manually checked for diversity (see SI Appendix, section SI 2.A for details). (3) We examined LLMs’ sensitivity to linguistic structure by introducing varying degrees of word order permutations to the query descriptions. We shuffled $30\%$, $60\%$ and $100\%$ of the words in the query descriptions and reinserted them into the original text.
@@
To probe the role of contextual information in concept inference, we examined how modifying the description–word pairings in context affected model predictions. For each target concept, we paired its description with a proxy symbol and combined it with N1N-1 correct description–word pairs from other concepts. We then queried the LLM using the same description. Model performance was evaluated based on how often the LLM replicated the proxy symbol or generated the correct word for the target concept. We tested four types of proxy symbols: (1) a random uppercase English letter, excluding “A” and “I” to avoid potential semantic associations; (2) a randomly generated lowercase letter string, with length sampled from a shifted Poisson distribution ($\textrm{mean}=6.94$, variance = 5.80) to approximate typical English word lengths [84]; (3) a random word selected from the THINGS database, distinct from the target concept and absent from the current context; and (4) the correct word for the target concept, used as a reference.
@@ 4.3 Structural analysis of representation spaces
We used representational similarity analysis (RSA) [42] to measure the alignment between conceptual representations derived under different conditions (i.e., across contexts or models), which is non-parametric and has been widely adopted to measure topological alignment between representation spaces [85]. Let $X\in\mathbb{R}^{m\times d_{1}}$ and $Y\in\mathbb{R}^{m\times d_{2}}$ denote two sets of conceptual representations for $m$ concepts with dimensionality of $d_{1}$ and $d_{2}$, respectively. Each space is characterized by its relational structure, represented by a (dis)similarity matrix $M\in\mathbb{R}^{m\times m}$, where the entry $M_{i,j}$ denotes the (dis)similarity between the representations of the $i^{\textrm{th}}$ and $j^{\textrm{th}}$ concepts in the corresponding space. The alignment can then be calculated as the Spearman’s rank correlation between the upper (or lower) diagonal portion of the two matrices $M_{X}$ and $M_{Y}$, yielding values ranging from 1-1 to $1$.
An appropriate similarity function is needed for the (dis)similarity matrix to characterize the relational structure of the representation space. While cosine similarity has been widely adopted since the advent of static word embeddings, it is sensitive to outliers and may be suboptimal for capturing non-linear relationships [86]. We thus employed two metrics: cosine similarity and Spearman’s rank correlation. Comparison with human behavioral data suggests that the cosine similarity captures semantic similarity while Spearman’s rank correlation reflects both similarity and association (SI Appendix, section SI 1.C; Fig. S11 AB). We adopted Spearman correlation as our primary similarity function, with results from cosine similarity provided in SI Appendix, section SI 1.B; Fig. S9 AB.
Additionally, we employed the parallelism score [58] to evaluate if the differences between conceptual representations are preserved across conditions, which have been thought to reflect relations in the representation space [46]. The results complemented the RSA findings (SI Appendix, section SI 1.B; Fig. S9 CD).
For qualitative comparison across LLMs, we visualized these models using t-SNE with a perplexity of 30 and 1000 iterations, where pairwise distances were defined as $1-\textrm{alignment}$.
4.4 Evaluation of model performance on downstream tasks
We evaluated LLMs’ language understanding and reasoning capabilities using a range of widely adopted multiple-choice benchmark tasks. Models were tested in a zero-shot setting by ranking candidate answers according to their assigned probabilities, with performance measured by accuracy in selecting the correct option. Full details of the benchmarks, prompt templates, and evaluation protocols are provided in SI Appendix, section SI 2.C; Table S3.
@@ 4.5 Characterization of model complexity
To characterize the complexity of the LLMs used in our experiments, we represented each by its number of parameters and amount of training data (i.e., the number of tokens used during training)—factors identified as crucial for determining model quality [87]. A principal component analysis (PCA) was then applied to these factors. The first principal component, accounting for $70.67\%$ of the total variance, was used to represent model complexity. For the Mistral models, where training data volume was not accessible, we estimated it based on comparable models released around the same time.
@@ 4.6 Prediction of human similarity judgments
To assess the alignment between LLM-derived representations and psychological measures of similarity, we first compared them with human ratings for concept pairs from SimLex-999 [49], available at https://fh295.github.io/simlex.html. SimLex-999 explicitly distinguishes semantic similarity from association or relatedness and contains human ratings for 999 concept pairs spanning 1,028 concepts across nouns, verbs and adjectives. We used description–word pairs from the WordNet data to provide contextual demonstrations to the LLMs, thereby deriving conceptual representations. Model performance was assessed through Spearman’s rank correlation between the similarity scores derived from the LLM representations and the human ratings.
We further employed the odd-one-out similarity judgments from the THINGS-data collection [39] to validate the effectiveness of LLM representations in handling the computation of similarities. The dataset consists of triplets sampled from the 1,854 concepts in THINGS, accessible at https://osf.io/f5rn6/. We presented LLMs with contextual demonstrations randomly sampled from THINGS to obtain their conceptual representations. For each triplet $\left(i,j,k\right)$, we computed pairwise similarities among the corresponding conceptual representations $\left(\mathbf{h}_{i},\mathbf{h}_{j},\mathbf{h}_{k}\right)$ and identified the item outside the most similar pair as the odd-one-out. Model predictions were compared with individual human judgments across the validation set (over 450,000 triplets).
@@ 4.7 Categorization
We used high-level human-labeled natural categories in THINGS [50]. We excluded subcategories of other categories, concepts belonging to multiple categories and categories with fewer than ten concepts [51]. This filtering resulted in 23 out of the original 53 categories, comprising 962 concepts in total (SI Appendix, Table S4). Categorization was evaluated using two approaches corresponding to prototype and exemplar models. For the prototype model, category membership was determined with a leave-one-out nearest-centroid classifier. In each iteration, the centroid for each category was computed by averaging the representations of the remaining concepts, and the left-out concept was assigned to the closest centroid. For the exemplar model, categorization followed a nearest-neighbor decision rule [88] (SI Appendix, section SI 2.E).
We combined multidimensional scaling (MDS) with t-SNE to visualize the representations in two dimensions, thereby preserving the global structure while better capturing local similarities. Representations were first reduced to 64 dimensions using MDS, with distances calculated as $1-\textrm{similarity}$, and then embedded in 2D using t-SNE with a perplexity of 30 and 1000 iterations.